OpenStudy (anonymous):

If y is a positive number, which of the following is equivalent to increasing y by 40 percent and then decreasing the result by 50%?

OpenStudy (anonymous):

\[(1.4)\times .5y\]

OpenStudy (anonymous):

where did you get 1.4 from?

OpenStudy (anonymous):

multiplying it by \(1.4\) increases it by \(40\%\) and \(.5\) decreases it by \(50\%\)

OpenStudy (anonymous):

for example, if you want to increase 30 by \(40\%\) you compute \(1.4\times 30\)

OpenStudy (anonymous):

I'm still not following :/ As a person with no prior knowledge of how to solve these mathematical problems, i don't see how you pull 1.4 out of the blue. How did you get 1.4 mathematically?

OpenStudy (anonymous):

100% plus 40% is 140%

OpenStudy (anonymous):

okay?

OpenStudy (anonymous):

suppose i want to increase 30 by 40% i would compute \(.4\times 30=12\) then add \(30+12=42\) the computation was \[30+.4\times 30=42\] by the distributive property \[30+.4\times 30=30(1+.4)=30\times 1.4\]

OpenStudy (anonymous):

What's the final answer?

OpenStudy (anonymous):

to increase a number by 20% multiply it by 1.2 to increase a number by 70% multiply it by .17 to decrease a number by 30% multiply it by .7 to decrease a number by 12% multiply it by .88

OpenStudy (anonymous):

whatever \(1.4\times .5\) is

OpenStudy (anonymous):

But that's wrong :/

OpenStudy (anonymous):

i assure you it is right

OpenStudy (anonymous):

I'm looking at my multiple choices and it's no where close to the answer you're telling me.

OpenStudy (anonymous):

@satellite73 Is correct.

OpenStudy (anonymous):

A number is originally at \(100\%\). If you decrease by \(30\%\) you get \(100\%-30\%=70\%\) or the original number.

OpenStudy (anonymous):

Percentages are found by dividing the percentage by \(100\) and multiplying that result to the number.

OpenStudy (anonymous):

\[ \frac {70}{100}=0.7 \]

OpenStudy (anonymous):

the multiple choices are: Increasing y by 20% , Increasing y by 10% , decreasing y by 10% , Decreasing y by 25% , Decreasing y by 30%

OpenStudy (anonymous):

You say 0.7 but it isn't listed in the multiple choice ! :(

OpenStudy (anonymous):

so increasing y by 10% ?

OpenStudy (anonymous):

\[ 70\%x=y \]

OpenStudy (anonymous):

\(70\%=100\%-30\%\)

OpenStudy (anonymous):

Decrease by \(30\%\)

OpenStudy (anonymous):

guess i owe satellite a huge apology!

OpenStudy (anonymous):

A huge apology to @Satellite73 I was being a huge brat. Thanks for trying to help me! I felt really bad. :(